OpenStudy (anonymous):

Cos(theta)= (1/3), find csc(90-(theta))

OpenStudy (debbieg):

Do the cofunction identities ring a bell? :) Use the cosine value to get the secant value. Then use the cofunction identity that relates cosine and cosecant... :)

OpenStudy (debbieg):

And, welcome to Open Study! :)

OpenStudy (anonymous):

first off \[\csc(90-\theta)=\sec(\theta)\]

OpenStudy (anonymous):

Since CSC(90-(theta) = Sec(theta). If I find the value of Theta for Secant will I get the right answer for CSC(90-(theta)?

OpenStudy (anonymous):

this is a "co function" identity

OpenStudy (anonymous):

secant is the reciprocal of cosine so all you need do is flip it

OpenStudy (anonymous):

Okay, I understand. Thank you.

OpenStudy (anonymous):

yw

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."